You wish to test the following hypotheses at a significance level of α=0.05α=0.05.

      H0:μ=62.1
      HA:μ<62.1

You believe the population is normally distributed, but you do not know the standard deviation. You obtain a sample of size n=90 with mean ¯x=60.7 and a standard deviation of s=7.6

What is the test statistic for this sample? (Report answer accurate to two decimal places.)


t=

What is the degree of freedom?

df=
Answer #1

Solution :

This is the left tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 62.1

Ha : \mu <  62.1

Test statistic = t

= (\bar x - \mu ) / s / \sqrt n

= (60.7 - 62.1) / 7.6 / \sqrt 90 =  -1.75

Test statistic = t = -1.75

sample size = n = 90

Degrees of freedom = df = n - 1 = 90 - 1 = 89